# Partial domain adaptation enables cross domain cell type annotation between scRNA-seq and snRNA-seq

### v2

@@ Abstract
Accurate cell type annotation across datasets is a key challenge in single-cell analysis. snRNA-seq enables profiling of frozen or difficult-to-dissociate tissues, complementing scRNA-seq by capturing fragile or rare cell types. However, cross-annotation between these two datasets remains largely unexplored, as existing methods treat them independently. We introduce ScNucAdapt, a method designed for cross-annotation between scRNA-seq and snRNA-seq datasets. To address distributional and cell composition differences, ScNucAdapt employs partial domain adaptation. Experiments across both unpaired and paired scRNA-seq and SnRNA-seq shows that ScNucAdapt achieves robust and accurate cell type annotation, outperforming existing approaches. Therefore, ScNucAdapt provides a practical framework for the cross-domain cell type annotation between scRNA-seq and snRNA-seq data.
@@ 1 Introduction
Cells are the fundamentals of life [regev2017human]. It is vital to annotate each cell correctly in terms of its transcriptomic profiles[pasquini2021automated], enabling the identification of distinct cellular populations, comparison across samples, and linkage of molecular profiles to biological function or disease[li2025overview]. Most published methods on automatic cell type annotations are based on ScRNA-seq, including Singlecellnet[tan2019singlecellnet], which uses an ensemble of Random Forest classifiers for the cell type annotation of ScRNA-seq datasets, and ScMap[kiselev2018scmap] works by comparing the gene-expression profile of each new cell to reference data and labeling the cell with the type that shows the highest similarity.
@@
However, when confronted with frozen samples or tissues that are difficult to dissociate, SnRNA-seq offers a practical alternative to scRNA-seq[ding2020systematic], capturing nuclear transcripts without viable whole cells and enabling detection of fragile or rare cell types that single-cell methods often underrepresent[wu2019advantages]. Many studies have integrated ScRNA-seq and SnRNA-seq, for instance, neurodegenerative diseases[zhang2023comprehensive], skeletal muscle[heuston2025optimized], frozen and fresh tumor samples[slyper2020single], and even PBMC for a disease progression study[park2025paired]. This shows that cross-domain annotation between scRNA-seq and snRNA-seq is essential for unifying cellular identities across two datasets and ensuring consistent interpretation of data generated from different tissue conditions or experimental protocols[quatredeniers2023meta]. Previous research on annotating cell types for snRNA-seq and scRNA-seq data relies on traditional machine learning methods[water2], particularly for kidney cell types[water]. However, these methods overlooked the relationships between scRNA-seq and snRNA-seq, treating them as separate datasets. Therefore, there’s an urgent need for development in cross-domain cell type annotation between scRNA-seq and snRNA-seq.
@@
However, Distributional differences often occur between scRNA-seq and SnRNA-seq[andrews2022single]. Moreover, in a real automatic annotation situation, the cell type composition of target datasets is unknown, which could cause cell type compositions to differ between the two datasets, making it challenging to achieve robust annotation between them. Therefore, inspired by partial domain adaptation, which can simultaneously address both the distribution differences between the two datasets and the mismatch in their label spaces, we developed a framework called ScNucAdapt that selectively transfers knowledge from the source dataset to the target dataset.
@@
Partial domain adaptation[li2022partial] addresses the problem of transferring knowledge from a labeled source domain to an unlabeled target domain when the target label space is a subset of the source label space, the concept is shown in Figure 1. Unlike traditional domain adaptation, which assumes identical label spaces across domains, Partial domain adaptation mitigates the negative transfer caused by irrelevant source classes. Partial domain adaptation has been applied to various fields, including fault diagnosis[zhu2023partial], pneumonia diagnosis from chest x-ray images[liu2023attention], and cross-session neural decoding[wang2025partial].
@@
This design enables ScNucAdapt to focus on cell types that are shared across datasets while minimizing the negative impact of non-overlapping or dataset-specific cell types. Moreover, ScJoint[scjoint], ScNCL[yan2023scncl], ScCobra[zhao2025sccobra], and ScCorrect[liu2025sccorrect], which utilize transfer learning and domain adaptation methods for label transfer from unpaired ScRNA-seq to ScATAC-seq datasets, also provided inspiration for developing ScNucAdapt.
@@
The following passages are organized as follows. In Section 2, the methods of ScNucAdapt are presented, and the descriptions of the datasets used and the evaluation metrics are included. In Section 3, the experimental results on classification accuracy between a set of ScRNA-seq and SnRNA-seq are presented, including an ablation experiment that demonstrates the effectiveness of each component in ScNucAdapt, as well as a sensitivity analysis. In Section 4, the discussion is presented. Finally, in Section 5, the conclusion is presented.
@@ 2.1 Shared Source and Target encoder
To extract features from both source and target datasets into a common label space, a shared encoder is used. The encoder is composed of two fully connected layers. The first layer transforms the input features into hidden units and applies a ReLU activation function to introduce non-linearity. The second layer reduces these features into a latent space, creating a compact representation that captures the most important patterns in the gene expression data.
2.2 Dynamic Selection of Clusters in Target data
In this section, we will introduce the concept of dynamic clustering in the target dataset[ronen2022deepdpm], without the prior knowledge of the number of clusters[huang2025praga]. The concept framework is presented in Figure 3.
Figure 3: Concept of Dynamic selection of Clusters in Target data
@@ 2.2 Dynamic Selection of Clusters in Target data
After obtaining the representation of the target dataset, we applied t-SNE[hinton2002stochastic] to the target representation for dimensionality reduction to reduce computational cost. Then, we set the initial cluster $C$ for the Gaussian mixture model to assign the target representation into $C$ clusters; note that $C$ doesn’t represent the true number of clusters in the target representation.
The mean and sample count of each cluster are denoted by $\mu_{c}$ and $N_{c}$, respectively, where the subscript $c$ refers to the cluster index. To enable dynamic adjustment of the total number of clusters, each cluster is further divided into two sub-clusters. The corresponding mean and sample count of the sub-clusters are represented as $\mu_{c,s}$ and $N_{c,s}$, where $s\in\{1,2\}$ denotes the sub-cluster index. A splitting criterion is then defined for each cluster to determine whether it should be further partitioned, as described in Eq.(2).
$S=\frac{\Gamma(N_{c,1})L(Z_{(c,1),\nu,\kappa})\Gamma(N_{c,2})L(Z_{(c,s)},\nu,\kappa)}{\Gamma(N_{c})L(Z_{c},\nu,\kappa)}$ (2)
where $\Gamma(\cdot)$ denotes the Gamma function, and $L(\cdot,\nu,\kappa)$ corresponds to the marginal likelihood evaluated under a Normal–Inverse–Wishart (NIW) prior, parameterized by the hyperparameters $\nu$ and $\kappa$. When $S>1$, the original cluster is substituted with one of its derived subclusters, and the remaining subcluster is incorporated as an additional, distinct cluster shown in Eq.(3).
$\mu_{c}\rightarrow\mu_{c,1};\mu_{C+1}\rightarrow\mu_{c,2}$ (3)
Moreover, we introduce a decision for merging, as shown in Eq. (4).
$M=\frac{\Gamma(N_{i}+N_{j})L(Z_{i}\cup Z_{j},\nu,\kappa)}{\Gamma(N_{i})L(Z_{i},\nu,\kappa)\Gamma(N_{j})L(Z_{j},\nu,\kappa)}$ (4)
If $M>1$, then the new merge cluster will replace the original two clusters with the average of the two clusters, shown in Eq. (5).
$\mu_{i}\rightarrow\varnothing;\mu_{j}\rightarrow\varnothing;\mu_{C-2+1}\rightarrow\frac{\mu_{i}+\mu_{j}}{2}$ (5)
@@ 2.3 Cauchy-Schwarz Divergence
In this section, we introduce Cauchy-Schwarz Divergence[yin2024domain]. Given the preprocessed source dataset $X_{s}\in R^{M\times m}$ and target dataset $X_{t}\in R^{N\times m}$, where the source data vectors $\{x_{s,i}\}\in X_{s}$ and target data vectors $\{x_{t,i}\}\in X_{t}$, the Cauchy–Schwarz Divergence (CS Divergence) is defined in Eq.(6).
$D_{CS}(X_{s};X_{t})=-\log{(\frac{(\int X_{s}(x_{s})X_{t}(x_{t})dz)^{2}}{\int X_{s}^{2}(x_{s})dz\int X_{t}^{2}(x_{t})dz})}$ (6)
where $X_{s}(x_{s})$ and $X_{t}(x_{t})$ are the kernel density estimators of source and target datasets, defined in Eq.(7) and Eq.(8).
$\hat{p}_{s}(x_{s})=\frac{1}{M}\sum^{M}_{i=1}\kappa_{\rho}(x_{s},x_{s,i})$ (7)
$\hat{p}_{t}(x_{t})=\frac{1}{N}\sum^{M}_{i=1}\kappa_{\rho}(x_{t},x_{t,i})$ (8)
Here we chose gaussian kernel function as the kernel estimator $\kappa_{\rho}(x,x^{{}^{\prime}})=\exp{(-\frac{\|x-x^{{}^{\prime}}\|^{2}_{2}}{2\sigma^{2}})}$. The calculation results of $\int\hat{p}_{s}^{2}(x_{s})$$\int\hat{p}_{t}^{2}(x_{t})$$\int\hat{p}_{s}(x_{s})\hat{p}_{t}(x_{t})$ are shown in Eq.(9), Eq.(10) and Eq.(11).
$\int\hat{p}_{s}^{2}(x_{s})=\frac{1}{M^{2}}\sum^{M}_{i=1}\sum^{M}_{j=1}\kappa_{\sigma}(x_{s,i},x_{s,j})$ (9)
$\int\hat{p}_{t}^{2}(x_{t})=\frac{1}{N^{2}}\sum^{N}_{i=1}\sum^{N}_{j=1}\kappa_{\sigma}(x_{t,i},x_{t,j})$ (10)
$\int\hat{p}_{s}(x_{s})\hat{p}_{t}(x_{t})=\frac{1}{MN}\sum^{M}_{i=1}\sum^{N}_{j=1}\kappa_{\sqrt{2}\sigma}(x_{t,j},x_{s,i})$ (11)
@@
$D_{CS}(X_{s};X_{t})$ $=\log(\frac{1}{M^{2}}\sum^{M}_{i,j=1}\kappa_{\sqrt{2}\sigma}(x_{s,j}-x_{s,i}))+$
$\log{(\frac{1}{N^{2}}\sum^{N}_{i,j=1}\kappa_{\sqrt{2}\sigma}(x_{t,j}-x_{t,i}))}$
$-2\log(\frac{1}{MN}\sum^{M}_{i=1}\sum^{N}_{i=1}\kappa_{\sqrt{2}\sigma}(x_{t,j},x_{s,i}))$
2.4 Decision for Merging
In this section, we will be introducing the decision for merging after obtaining the CS divergence between source known cell clusters and predicted target clusters.
Given the source dataset $X_{s}\in R^{m\times n}$ with $p$ labels, we could divide the source dataset into $p$ subsets, noting as $X_{s}=\{X_{1,s},X_{2,s},\dots,X_{p,s}\}$. After performing dynamic clustering on the target datasets, we obtain $\hat{C}$ clusters $X_{t}=\{X_{1,t},X_{2,t},\dots,X_{\hat{C},t}\}$. Then, in terms of CS divergence described in Section 2.3, we calculate the CS divergence of each pair from the source and target subsets, $CS_{i,j}(X_{i,s},X_{j,t})$, where $i=1,2,\dots,p$ and $j=1,2,\dots,\hat{C}$. To identify which subsets in the source datasets best correspond to each cluster in the target datasets, we selected those with the lowest CS divergence paired with each target cluster. Therefore, we could obtain the total loss function for minimization shown in Eq.(12), where $a_{i}$ represents the corresponding subsets that match the i-th target cluster.
$L_{cs}=\sum^{\hat{C}}_{i=1}D_{CS}(X_{a_{i},s},X_{i,t})$ (12)

### v3

@@ Abstract
Accurate cell type annotation across datasets is a key challenge in single-cell analysis. snRNA-seq enables profiling of frozen or difficult-to-dissociate tissues, complementing scRNA-seq by capturing fragile or rare cell types. However, cross-annotation between these two datasets remains largely unexplored, as existing methods treat them independently. We introduce ScNucAdapt, a method designed for cross-annotation between paired and unpaired scRNA-seq and snRNA-seq datasets. To address distributional and cell composition differences, ScNucAdapt employs partial domain adaptation. Experiments across both unpaired and paired scRNA-seq and snRNA-seq show that ScNucAdapt achieves robust and accurate cell type annotation, outperforming existing approaches. Therefore, ScNucAdapt provides a practical framework for the cross-domain cell type annotation between scRNA-seq and snRNA-seq data.
@@
Author Summary
Single-cell and single-nucleus RNA sequencing are two powerful technologies that allow scientists to study gene activity in individual cells. However, comparing data between these methods remains challenging because they capture different parts of the cell and are often collected under different conditions. This makes it difficult to consistently identify cell types across experiments, hindering our understanding of health and disease.
We developed ScNucAdapt, a computational framework that can automatically transfer cell type knowledge between these two types of datasets, even when they come from different laboratories or tissue conditions. Our method learns to recognize shared patterns while ignoring dataset-specific differences. Through testing on diverse tissues, including bladder, kidney, tumors, and brain, we show that ScNucAdapt consistently outperforms existing approaches.
By enabling reliable integration of single-cell and single-nucleus data, our work helps researchers build more complete pictures of cellular diversity across tissues and disease states. This capability is particularly valuable for studying archived frozen samples or fragile cell types that are difficult to analyze with conventional methods, potentially accelerating discoveries in various fields.
@@ 1 Introduction
Cells are the fundamentals of life [23]. It is vital to annotate each cell correctly in terms of its transcriptomic profiles [21], enabling the identification of distinct cellular populations, comparison across samples, and linkage of molecular profiles to biological function or disease [12]. Most published methods on automatic cell type annotations are based on scRNA-seq, including SingleCellNet [28], which uses an ensemble of Random Forest classifiers for the cell type annotation of scRNA-seq datasets, and ScMap [9] works by comparing the gene-expression profile of each new cell to reference data and labeling the cell with the type that shows the highest similarity.
@@
However, when confronted with frozen samples or tissues that are difficult to dissociate, snRNA-seq offers a practical alternative to scRNA-seq [4], capturing nuclear transcripts without viable whole cells and enabling detection of fragile or rare cell types that single-cell methods often underrepresent [33]. Many studies have integrated scRNA-seq and snRNA-seq, for instance, neurodegenerative diseases [38], skeletal muscle [7], frozen and fresh tumor samples [27], and even PBMC for a disease progression study [20]. This shows that cross-domain annotation between scRNA-seq and snRNA-seq is essential for unifying cellular identities across two datasets and ensuring consistent interpretation of data generated from different tissue conditions or experimental protocols [22]. Previous research on annotating cell types for snRNA-seq and scRNA-seq data relies on traditional machine learning methods [11], particularly for kidney cell types [29]. However, these methods overlooked the relationships between scRNA-seq and snRNA-seq, treating them as separate datasets. Therefore, there’s an urgent need for development in cross-domain cell type annotation between scRNA-seq and snRNA-seq.
Figure 1: Concept of pair and unpaired scRNA-seq and snRNA-seq
@@
However, Distributional differences often occur between scRNA-seq and snRNA-seq [1], including paired and unpaired ones, as shown in Figure 1. Moreover, in a real automatic annotation situation, the cell type composition of target datasets is unknown, which could cause cell type compositions to differ between the two datasets, making it challenging to achieve robust annotation between them. Therefore, inspired by partial domain adaptation, which can simultaneously address both the distribution differences between the two datasets and the mismatch in their label spaces, we developed a framework called ScNucAdapt that selectively transfers knowledge from the source dataset to the target dataset.
@@
Partial domain adaptation [13] addresses the problem of transferring knowledge from a labeled source domain to an unlabeled target domain when the target label space is a subset of the source label space, as shown in Figure 2. Unlike traditional domain adaptation, which assumes identical label spaces across domains, Partial domain adaptation mitigates the negative transfer caused by irrelevant source classes. Partial domain adaptation has been applied to various fields, including fault diagnosis [41], pneumonia diagnosis from chest x-ray images [16], and cross-session neural decoding [30].
@@
This design enables ScNucAdapt to focus on cell types that are shared across datasets while minimizing the negative impact of non-overlapping or dataset-specific cell types. Moreover, ScJoint [14], ScNCL [34], ScCobra [39], and ScCorrect [17], which utilize transfer learning and domain adaptation methods for label transfer from unpaired scRNA-seq to ScATAC-seq datasets, also provided inspiration for developing ScNucAdapt.
@@
The following passages are organized as follows. In Section 2, the methods of ScNucAdapt are presented, and the descriptions of the datasets used and the evaluation metrics are included. In Section 3, the experimental results on classification accuracy between a set of scRNA-seq and snRNA-seq are presented, including an ablation experiment that demonstrates the effectiveness of each component in ScNucAdapt, as well as a sensitivity analysis and a runtime and memory scaling test. In Section 4, the discussion is presented. Finally, in Section 5, the conclusion is presented.
@@ 2.1 Shared Source and Target encoder
To extract features from both source and target datasets into a common label space, a shared encoder is used. The encoder is composed of two fully connected layers. The first layer transforms the input features into hidden units. The second layer reduces these features into a latent space, creating a compact representation that captures the most important patterns in the gene expression data.
2.2 Dynamic clustering in Target data
In this section, we will introduce the concept of dynamic clustering in the target dataset without the prior knowledge of the number of clusters, which was inspired by DeepDPM [24] for an unknown number of clusters for deep clustering, and PRAGA [8] for spatial multi-omics clustering.
@@ 2.2 Dynamic clustering in Target data
After obtaining the representations of the target dataset, we set the initial cluster $C$ for the Gaussian mixture model to assign the target representations into $C$ clusters; note that $C$ doesn’t represent the true number of clusters in the target representations or the true cell type labels in the source dataset, and would further be adjusted through a split and merge framework, which is performed through the Metropolis-Hastings framework [5].
The sample count and the target representations of each cluster are denoted by $N_{c}$ and $Z^{t}_{c}$, respectively, where the subscript $c=1,2,\dots,C$ denotes the cluster index. To enable dynamic adjustment of the total number of clusters, each cluster is further divided into two sub-clusters using Gaussian mixture. The corresponding sample count of the sub-clusters is represented as $N_{c,s}$, where $s\in\{1,2\}$ denotes the sub-cluster index. A splitting criterion is then defined for each cluster to determine whether it should be further partitioned, where the split is accepted with probability $\min{(1,H_{s})}$. The hasting ratio $H_{s}$ is defined in Eq. (2).
$H_{s}=\frac{\Gamma(N_{c,1})L(Z^{t}_{c,1};\nu,\kappa,m,\psi)\Gamma(N_{c,2})L(Z^{t}_{c,2};\nu,\kappa,m,\psi)}{\Gamma(N_{c})L(Z^{t}_{c};\nu,\kappa,m,\psi)}$ (2)
where $\Gamma(\cdot)$ denotes the Gamma function, and $L(\cdot;\nu,\kappa,m,\psi)$ corresponds to the marginal likelihood evaluated under a Normal–Inverse–Wishart (NIW) prior, parameterized by the hyperparameters $\nu$, $\kappa$, $m$, $\psi$. When $H_{s}>1$, the original cluster is substituted with one of its derived subclusters, and the remaining subcluster is incorporated as an additional, distinct cluster shown in Eq. (3).
$Z^{t}_{c}:=Z^{t}_{c,1};Z^{t}_{C}:=Z^{t}_{c,2}(C:=C+1)$ (3)
After splitting the clusters, we introduce a decision for merging criterion based on the clusters after splitting, where the merging is accepted with probability $\min{(1,H_{m})}$, as shown in Eq. (4).
$H_{m(i,j)}=\frac{1}{H_{s(i,j)}}=\frac{\Gamma(N_{i}+N_{j})L(Z^{t}_{i}\cup Z^{t}_{j};\nu,\kappa,m,\psi)}{\Gamma(N_{i})L(Z^{t}_{i};\nu,\kappa,m,\psi)\Gamma(N_{j})L(Z^{t}_{j};\nu,\kappa,m,\psi)}$ (4)
Rather than exhaustively considering all possible merges in sequence, we limit the merge candidates for each cluster $N_{i}$ to its nearest neighbors $N_{j}$. If $H_{m(i,j)}>1$, then the new merge cluster will replace the original two clusters, shown in Eq. (5).
$Z^{t}_{i}:=\varnothing;Z^{t}_{j}:=\varnothing;Z^{t}_{C}:=Z^{t}_{i}\cup Z^{t}_{j}(C:=C-1)$ (5)
@@ 2.3 Cauchy-Schwarz Divergence
In this section, we introduce Cauchy-Schwarz Divergence [35] and the emperical estimator of $D_{CS}(p_{s}(z);p_{t}(z))$. Given the source representations of a certain class $Z_{s}\in R^{N\times m^{{}^{\prime}}}$ and target representations of a certain cluster $Z_{t}\in R^{n\times m^{{}^{\prime}}}$, where the source representations vectors $\{z_{i}^{s}\}_{i=1}^{N}\in Z_{s}$ and target representations vectors $\{z_{i}^{t}\}_{i=1}^{n}\in Z_{t}$, the Cauchy-Schwarz Divergence (CS Divergence) is defined in Eq. (6).
$D_{CS}(p_{s};p_{t})=-\log{(\frac{(\int p_{s}(z)p_{t}(z)\mathrm{d}z)^{2}}{\int p_{s}^{2}(z)\mathrm{d}z\int p_{t}^{2}(z)\mathrm{d}z})}$ (6)
Using kernel density estimation, we estimate the densities from finite samples defined in Eq. (7) and Eq. (8).
$\hat{p}_{s}(z)=\frac{1}{M}\sum^{M}_{i=1}\kappa_{\sigma}(z,z_{i}^{s})$ (7)
$\hat{p}_{t}(z)=\frac{1}{N}\sum^{M}_{i=1}\kappa_{\sigma}(z,z_{i}^{t})$ (8)
Here we chose gaussian kernel function as the kernel estimator $\kappa_{\sigma}(z,z^{{}^{\prime}})=\exp{(-\frac{\|z-z^{{}^{\prime}}\|^{2}_{2}}{2\sigma^{2}})}$, where $\sigma$ is the bandwidth parameter that controls the smoothness of the kernel. The calculation results of $\int\hat{p}_{s}^{2}(z)\mathrm{d}z$, $\int\hat{p}_{t}^{2}(z)\mathrm{d}z$, $\int\hat{p}_{s}(z)\hat{p}_{t}(z)\mathrm{d}z$ are shown in Eq. (9), Eq. (10) and Eq. (11).
$\int\hat{p}_{s}^{2}(z)\mathrm{d}z=\frac{1}{M^{2}}\sum^{M}_{i=1}\sum^{M}_{j=1}\kappa_{\sqrt{2}\sigma}(z_{j}^{s},z_{i}^{s})$ (9)
$\int\hat{p}_{t}^{2}(z)\mathrm{d}z=\frac{1}{N^{2}}\sum^{N}_{i=1}\sum^{N}_{j=1}\kappa_{\sqrt{2}\sigma}(z_{j}^{t},x_{i}^{t})$ (10)
$\int\hat{p}_{s}(x_{s})\hat{p}_{t}(x_{t})\mathrm{d}z=\frac{1}{MN}\sum^{M}_{i=1}\sum^{N}_{j=1}\kappa_{\sqrt{2}\sigma}(z_{j}^{t},z_{i}^{t})$ (11)
@@
$D_{CS}(p_{s};p_{t})$ $=\log(\frac{1}{M^{2}}\sum^{M}_{i,j=1}\kappa_{\sqrt{2}\sigma}(z_{j}^{s}-z_{i}^{s}))+$
$\log{(\frac{1}{N^{2}}\sum^{N}_{i,j=1}\kappa_{\sqrt{2}\sigma}(z_{j}^{t}-z_{i}^{t}))}$
$-2\log(\frac{1}{MN}\sum^{M}_{i=1}\sum^{N}_{j=1}\kappa_{\sqrt{2}\sigma}(z_{j}^{t},z_{i}^{s}))$
2.4 Source Class-Target Cluster Matching
In this section, we will be introducing the Source Class-Target Cluster Matching after obtaining the CS divergence between source known cell classes and predicted target clusters. And further assign the matched pairs’ CS Divergence to the final training loss.
Given the source representations vectors $Z_{s}$ and with $p$ labels, we could divide the source representations into $p$ subsets, noting as $Z_{s}=\{Z_{1,s},Z_{2,s},\dots,Z_{p,s}\}$. After performing dynamic clustering on the target representations on target representations vectors $Z_{t}$, we obtain $\hat{C}$ clusters $Z_{t}=\{Z_{1,t},Z_{2,t},\dots,Z_{\hat{C},t}\}$. Then, in terms of CS divergence described in Section 2.3, we calculate the CS divergence of each pair from the source classes and target clusters $Z_{i,s}$ and $Z_{j,t}$, which is $D_{CS}(p_{i,s};p_{j,t})$, where $i=1,2,\dots,p$ and $j=1,2,\dots,\hat{C}$. To identify which subsets in the source representations best correspond to each cluster in the target representations, we selected those with the lowest CS divergence paired with each target cluster. Therefore, we could obtain the global merge decision, the total loss function for minimization on the mini-batched source and target dataset are shown in Eq. (12), where $a_{i}$ represents the corresponding subsets that match the i-th target cluster.
$L_{cs}=\sum^{\hat{C}}_{i=1}D_{CS}(p_{a_{i},s}^{b};p_{i,t}^{b})$ (12)